# CrystalFormer-RL: Reinforcement Fine-Tuning for Materials Design

### v1

@@ Abstract
Reinforcement fine-tuning has instrumental enhanced the instruction-following and reasoning abilities of large language models. In this work, we explore the applications of reinforcement fine-tuning to the autoregressive transformer-based materials generative model CrystalFormer (arXiv:2403.15734) using discriminative machine learning models such as interatomic potentials and property prediction models. By optimizing reward signals—such as energy above the convex hull and material property figures of merit—reinforcement fine-tuning infuses knowledge from discriminative models into generative models. The resulting model, CrystalFormer-RL, shows enhanced stability in generated crystals and successfully discovers crystals with desirable yet conflicting material properties, such as substantial dielectric constant and band gap simultaneously. Notably, we observe that reinforcement fine-tuning enables not only the property-guided novel material design ability of generative pre-trained model but also unlocks property-driven material retrieval from the unsupervised pre-training dataset. Leveraging rewards from discriminative models to fine-tune materials generative models opens an exciting gateway to the synergies of the machine learning ecosystem for materials.
@@ I Introduction
In materials science, machine learning (ML) techniques are revolutionizing research to enable rapid design and discovery of materials with desired properties. Machine learning models can be broadly categorized into generative and discriminative approaches. Crystal generative models, aim to capture the prior probability distribution $p(x)$ of stable crystal structures in the chemical space, learning the underlying patterns and correlations in crystal data to generate novel structures. These generative models [1, 2, 3, 4, 5, 6] capture the underlying distribution of stable materials in the chemical space and enable the direct generation of novel materials. On the other hand, discriminative models, including machine learning interatomic potentials (MLIP) [7, 8, 9, 10, 11, 12, 13] and property prediction models [14, 15], focus on modeling the conditional probability $p(y|x)$, where $y$ represents material properties given a crystal structure $x$. These models excel at predicting specific properties or behaviors of materials and enable accelerated simulations and property predictions. The integration of these complementary approaches—combining the generative capability of models that capture $p(x)$ with the predictive power of discriminative models that capture $p(y|x)$—presents a powerful framework for materials discovery. Further achievements hinge on the flexible combination of discriminative and generative models, enhancing their performance and adapting them to specific application domains. Reinforcement learning (RL) provides a principled way to achieve this synergy, allowing the generative model to be improved and guided by the knowledge embedded in discriminative models.
@@
The synergy of discriminative and generative machine learning models has already shown great success in real-world applications in the past few years. A notable example is classifier-guided generation [16, 17], where supervised trained discriminative models are leveraged to guide sampling of generative models, eliminating the need for iterative retraining of the generative model. Moreover, in the post-training phase of large language models (LLM), reinforcement learning from human feedback (RLHF) [18, 19] is employed to extract signals from supervised trained reward models to instruct generative models. In this context, RL [20] serves as a powerful mechanism to integrate unsupervised and supervised learned models. For example, LLM trained via RL have demonstrated improved instruction following [18] and reasoning capabilities [21], which shows it is an effective way to enhance the performance of generative models.
@@
Inspired by RLHF, we devise an RL approach to finetune the crystal generative model CrystalFormer using discriminative models such as MLIP and property prediction models, see Fig. 1(a). We demonstrate that the approach effectively enhances the stability and desired figures of merit of generated materials. We have released the codes and trained models at [22]. Moreover, we provide instructions on fine-tuning CrystalFormer with customized MLIP or property prediction models other than the ones used in this work.
@@ II Method
$\mathcal{L}=\mathop{\mathbb{E}}_{x\sim p_{\theta}(x)}\left[r(x)-\tau\ln\frac{p_{\theta}(x)}{p_{\mathrm{base}}(x)}\right],$ (1)
@@
The combination of both terms in Eq. (1) can also be regarded as a KL divergence between the policy network $p_{\theta}(x)$ under fine-tuning and an optimal policy [23, 25]
@@
where $Z$ is the normalization factor. The corresponding optimal value of the objective function Eq. (1) will be $\tau\ln Z$. Setting the reward to be proportional to the log-likelihood of certain desired property $r(x)\propto\tau\ln p(y|x)$, one sees that $p^{*}_{\theta}(x)\propto p_{\mathrm{base}}(x)p(y|x)$. Therefore, Eq. (2) can be viewed as Bayesian inference with a prior distribution $p_{\mathrm{base}}(x)$ and the likelihood $\exp\left(r(x)/{\tau}\right)$ [26, 25]. The pre-trained model not only provides the prior distribution $p_{\mathrm{base}}(x)$ but also provides parameter initialization for the policy network. Figure 1(b) illustrates the RL fine-tuning process. The variational inference provides an alternative to the Markov chain Monte Carlo (MCMC) sampling carried out in Ref. [5]. Here, larger $\tau$ corresponds to higher temperature and hence stronger entropy regularization. Such optimization is also equivalent to the variational free energy calculations carried out for statistical mechanics problems [27, 28], where the two terms in the objective function Eq. (1) correspond to expected energy and entropy, respectively. We employ the proximal policy optimization (PPO) [29] algorithm to maximize the expected reward. The implementation details are provided in Appendix C.
@@
The CrystalFormer is an autoregressive transformer which models the tokenized representation of crystal structures with explicit knowledge of space group symmetries. The representation includes the space group number, Wyckoff letter, chemical element, and fractional coordinates of each symmetry inequivalent atom, and finally, the lattice parameters of the unit cell [5]. For example, the crystal LaH10 in the $Fm\bar{3}m$ (No. 225) space group with the cubic conventional unit cell of the length $5.1\AA$ is represented as the sequence ”225-a-La-0-0-0-c-H-1/4-1/4-1/4-f-H-0.375-0.375-0.375-X-5.1-5.1-5.1-90-90-90”, which contains all the necessary information of the crystal. Key to the design of CrystalFormer is the sequential nature of Wyckoff letters which decrease in site symmetry in an alphabet order. In this regard, the CrystalFormer leverages Nature’s codebook–the Wyckoff position table–for a multimodal probabilistic modeling. Although the representation appears to lack a sense of spatial geometry, unsupervised pre-training allows the model to ingest solid-state chemistry knowledge by compressing the crystalline material database into the model parameters. Furthermore, the RL phase may allow it to extract the geometry and physics from MLIP and property prediction models trained with supervised approaches.
@@ III.1 Reinforcement learning from MLIP for improved stability of generated materials
As the first application of the reinforcement fine-tuning of CrystalFormer, we choose the energy above the convex hull as the reward signal, i.e., $r(x)=-E_{\mathrm{hull}}(x)$. The energy above the convex hull is the key metric for assessing the stability of a material, with lower values indicating greater stability. To predict the energy and calculate the energy above the convex hull based on the Alexandria convex hull [37], we utilize the Orb model [13]. This choice allows us to obtain accurate energy predictions and assess the stability of materials relative to the convex hull fast. It is important to emphasize that our approach is not limited to the Orb model; any universal MLIP can serve as the reward model. More accurate MLIP yields better reward signals, enhancing alignment with DFT calculations. Benchmarks on universal MLIP such as [38] provide an interactive leaderboard to select the favorable one for the fine-tuning purpose. We evaluate the performance by measuring the fraction of structures that are stable, unique, and novel (S.U.N.) [3]. The stability is defined as the energy above the convex hull being less than 0.1 eV/atom, uniqueness is defined as the generated structure not being identical to any other generated structures, and novelty is defined as the generated structure not being present in the training dataset.
@@
Figure 2(a) shows the expected energy above the convex hull of generated samples, which shows steady improvement during the training process. Figure 2(b) shows the KL divergence between the policy and the pre-trained base model $\mathrm{KL}(p_{\theta}\|p_{\mathrm{base}})=\mathop{\mathbb{E}}_{x\sim p_{\theta}(x)}\left[\ln\frac{p_{\theta}(x)}{p_{\mathrm{base}}(x)}\right]$, which slightly increases during the fine-tuning process.
@@
Figure 3: The histogram of energy above convex hull predicted by the Orb model [13] for the crystal samples from the pre-trained and the RL fine-tuned model respectively. The red dashed line indicates the threshold of 0.1 eV/atom. RL finetune improves the ratio of stable materials from 44.7% in the base model to 73.4%.
Figure 4: Fraction of generated structures that are stable, and are also S.U.N. for 14 space groups spanning the seven crystal systems.
@@
Figure 3 shows the histogram of the energy above the convex hull for the pre-trained model and the reinforcement fine-tuned model. For each model, 1000 structures were sampled, with their space groups drawn from the training data distribution. We note that we do not relax the generated structures, as the focus of the fine-tuning is to enhance the stability of generated materials. For the pre-trained model, 44.7% of the generated structures are stable, whereas RL improves this ratio to 73.4%, representing a substantial improvement. Simultaneously, the space group symmetry controlled  [^1] We sample space group index following the same distribution of the training dataset and use it as precondition to the CrystalFormer. S.U.N. ratio across all generated structures increased from 15.3% for the pre-trained model to 21.6% for the RL fine-tuned model. The results indicate that the RL fine-tuned model generates structures with lower energy above the convex hull compared to the pre-trained base model, suggesting that the RL method effectively optimizes the model to generate more stable structures. Besides reducing the energy above the convex hull, the reinforcement fine-tuning also improves the structure validity metrics compared to the base model [1], see Appendix B.
@@
Figure 4 shows the fraction of generated structures that are stable and are also S.U.N. for 14 space groups spanning the seven crystal systems. We choose to investigate the performance of the model on the same 14 space groups as Ref. [3]. For each space group, we generate 1000 structures and evaluate the quality of samples using the Orb model [13]. The results demonstrate that the reinforcement fine-tuned model generates a higher fraction of stable and S.U.N. structures in most cases, indicating that the RL method can enhance the model’s ability to generate S.U.N. structures even though we only award for the stability function. For space groups with high symmetries (e.g. $F\bar{4}3m$), even though the fraction of stable materials increases, the S.U.N. ratio does not increase possibly due to the restricted degrees of freedom limiting the structural motifs. From the figure one also observes that the performance of CrystalFormer is worst for the $P1$ space group where there is minimal symmetry-related inductive bias to be explored. This is in sharp contrast to Ref. [3] which shows the best S.U.N. in the $P1$ space group. In fact, for space group with higher symmetry which are more relevant to the natural distribution of inorganic materials [40], we observed that the performance of CrystalFormer is on par or even better than the ones shown in Fig. D8 of Ref. [3]. Lastly, while reinforcement fine-tuning does not affect the generation speed of the model, it increases the rate of generating S.U.N. samples from 3.46 to 4.89 per second.
@@
Figure 5 breaks down the performance enhancement in terms of the S.U.N. ratio coming from enlarging the pre-training dataset and different ways of fine-tuning. First of all, enlarging the pre-training dataset from MP-20 to Alex-20 yields a significant improvement in the S.U.N. ratio, rising from 2.6% to 15.3%. Next, we also tested the improvement brought by supervised fine-tuning (SFT) on the pre-trained model. For SFT, we generate 10,000 crystal samples and relax them using the Orb models. Only those stable samples with $E_{\mathrm{hull}}<0.1$ eV/atom were used to finetune the pre-trained model. Figure 5 shows that SFT yields some improvement over the pre-trained model, further increasing the S.U.N. ratio from 15.3% to 18.3%. However, it is the reinforcement fine-tuned model that delivers the most significant increase in the S.U.N. ratio, leading to a final S.U.N. ratio of 21.6%.
@@ III.2 Reinforcement learning from the figure of merits for property-guided material generation
Figure 6: (a) ML-predicted band gap $E_{g}$ versus total dielectric constant $\varepsilon_{\mathrm{total}}$ for generated samples in the $Fm\bar{3}m$ space group (No.225). The dashed line indicates a constant figure of merit at values $\varepsilon_{\mathrm{total}}\cdot E_{g}=100$eV. Note that we subsample 1000 data points from the generated samples for better visualization. (b) The histogram of energy above the convex hull predicted by the Orb model [13] for the crystal samples from the pre-trained and the RL fine-tuned model in $Fm\bar{3}m$ space group (No.225), respectively. The red dashed line indicates the threshold of 0.1 eV/atom. RL improves the FoM of the generated samples towards high FoM dielectric materials at the cost of only slightly affecting the stability. (c) Two examples of discovered high FoM materials along with their properties and tokenized representation (fractional coordinate and lattice parameters omitted).
Next, we apply the RL method for dielectric material discovery. Dielectric materials are essential components in a wide range of contemporary electronic devices, including flash storage, central processing units, and random access memory [41, 42]. The functional utility of these materials is fundamentally governed by the precise interplay between the dielectric constant and band gap—two inherently anti-correlated properties that seldom achieve optimality concurrently in a single material. Consequently, delineating a Pareto front becomes essential for navigating the trade-offs between these competing parameters. We utilize the pre-trained MEGNet [15, 43] model to predict the band gap $E_{g}$ for generated samples. The total dielectric constant is a dimensionless number, which is the ratio of the permittivity of a substance to the permittivity of vacuum. To predict the total dielectric constant $\varepsilon_{\mathrm{total}}$ which includes the ionic and electronic contributions [44], we train a separate M3GNet [30] model on the dataset of density functional perturbation theory (DFPT) calculations in the Materials Projects [45, 46], more details of training and evaluation can be found in Appendix D. The reward function is defined as the figure of merit (FoM), $r(x)=\varepsilon_{\mathrm{total}}\cdot E_{g}$, which favors materials with both sizeable band gap and permittivity.
@@
Figure 6(a) shows the ML predicted band gap $E_{g}$ versus total dielectric constant $\varepsilon_{\mathrm{total}}$ for generated samples in the $Fm\bar{3}m$ space group (No.225) using the base and fine-tuned model respectively. The reinforcement fine-tuning significantly improves the FoM, with the majority of generated samples exhibiting higher FoM values compared to the pre-trained base model. We observe that the RL method primarily enhances the FoM by increasing the band gap, rather than the dielectric constant. This outcome may be attributed to the limited and relatively low dielectric constant data used to train the dielectric model, potentially causing the model to predict lower dielectric constants. To address this issue, a more robust dielectric prediction model that captures a wider range of dielectric constants could be employed [47]. In contrast, the band gap prediction model demonstrates greater accuracy, resulting in the RL method favoring an increase in the band gap to optimize the FoM.
@@
We generated 5000 samples (which takes 276 seconds with single A100 GPU) and list those with $E_{\mathrm{hull}}<0.1$ eV/atom and $\mathrm{FoM}>100$eV in Table LABEL:table:dielectric_samples in Appendix D. We observed that many of these samples are fluoride, probably due to fluorine’s strong electronegativity increases the ionic character of bonds, so that to ensure a sizeable band gap. Moreover, many of these high FoM samples are composed of Cs elements, which may be due to the relatively large ionic radius of Cesium, leading to higher polarizability and thus an increased total dielectric constant of the material. Figure 6(b) shows the energy above the convex hull for the materials sampled from the base and fine-tuned model respectively, which suggests that the reinforcement fine-tuning does not significantly affect the stability of generated materials. Two examples of high FoM materials discovered by CrystalFormer-RL, Cs2BaSrF6 and Li6PbCl8, are shown in Fig 6(c). Cs2BaSrF6 belongs to the double perovskite [48] structure prototype. We note that although it is flagged as novel in Table LABEL:table:dielectric_samples, it is related to partial substitution [49] of the Ba element in CsBaF3 which is in the Alex-20 dataset. Therefore, the synthesis and identification of Cs2BaSrF6 as a novel compound may subject to caveats pointed out in Ref. [50] such as compositional disorder. Li6PbCl8 is derived from Li6VCl8 [51] by substituting V with Pb and belongs to the Suzuki structure prototype [52].

### v2

@@ Abstract
Reinforcement fine-tuning played an instrumental role in enhancing the instruction-following and reasoning abilities of large language models. In this work, we employ reinforcement fine-tuning for materials design, in which discriminative machine learning models are used to provide rewards to the autoregressive transformer-based materials generative model CrystalFormer [1]. By optimizing the reward signals—such as energy above the convex hull and material property figures of merit—reinforcement fine-tuning infuses knowledge from discriminative models into generative models. The resulting model, CrystalFormer-RL, shows enhanced stability in generated crystals and successfully discovers crystals with desirable yet conflicting material properties, such as substantial dielectric constant and band gap simultaneously. Notably, we observe that reinforcement fine-tuning not only enables the property-guided material design but also unlocks property-based material retrieval behavior of pretrained generative model. The present framework opens an exciting gateway to the synergies of the machine learning ecosystem for materials design.
@@ I Introduction
In materials science, machine learning (ML) techniques are revolutionizing research to enable rapid design and discovery of materials with desired properties. Machine learning models can be broadly categorized into generative and discriminative approaches. Crystal generative models, aim to capture the prior probability distribution $p(x)$ of stable crystal structures in the chemical space, learning the underlying patterns and correlations in crystal data to generate novel structures. These generative models [2, 3, 4, 5, 1, 6] capture the underlying distribution of stable materials in the chemical space and enable the direct generation of novel materials. On the other hand, discriminative models, including machine learning interatomic potentials (MLIP) [7, 8, 9, 10, 11, 12, 13] and property prediction models [14, 15], focus on modeling the conditional probability $p(y|x)$, where $y$ represents material properties given a crystal structure $x$. These models excel at predicting specific properties or behaviors of materials and enable accelerated simulations and property predictions. The integration of these complementary approaches—combining the generative capability of models that capture $p(x)$ with the predictive power of discriminative models that capture p(y|x)p(y|x)—presents a powerful framework for materials discovery. Further achievements hinge on the flexible combination of discriminative and generative models, enhancing their performance and adapting them to specific application domains. Reinforcement learning (RL) provides a principled way to achieve this synergy, allowing the generative model to be improved and guided by the knowledge embedded in discriminative models.
@@
The synergy of discriminative and generative machine learning models has already shown great success in real-world applications in the past few years. A notable example is classifier-guided generation [16, 17], where supervised trained discriminative models are leveraged to guide sampling of generative models, eliminating the need for repeated retraining of the generative model. Moreover, in the post-training phase of large language models (LLM), reinforcement learning from human feedback (RLHF) [18, 19] is employed to extract signals from supervised trained reward models to instruct generative models. In this context, RL [20] serves as a powerful mechanism to integrate unsupervised and supervised learned models. For example, LLM trained via RL have demonstrated improved instruction following [18] and reasoning capabilities [21], which shows it is an effective way to enhance the performance of generative models.
@@
Inspired by RLHF, we devise an RL approach to finetune the crystal generative model CrystalFormer using the reward provided by discriminative models such as MLIP and property prediction models, see Fig. 1(a). We demonstrate that the approach effectively enhances the stability and desired figures of merit of generated materials. We have released the codes and trained models [22]. Moreover, we provide instructions on fine-tuning CrystalFormer with customized MLIP or property prediction models for broader applications.
@@ II Method
$\mathcal{L}=\mathop{\mathbb{E}}_{x\sim p_{\theta}(x)}\left[r(x)-\tau\ln\left(\frac{p_{\theta}(x)}{p_{\mathrm{base}}(x)}\right)\right],$ (1)
@@
The combination of both terms in Eq. (1) can also be regarded as a KL divergence between the policy network $p_{\theta}(x)$ under fine-tuning and the optimal policy [23, 25]
@@
where $Z$ is the normalization factor. The corresponding optimal value of the objective function Eq. (1) will be $\tau\ln Z$. Setting the reward to be proportional to the log-likelihood of certain desired property $r(x)\propto\tau\ln p(y|x)$, one sees that $p^{*}_{\theta}(x)\propto p_{\mathrm{base}}(x)p(y|x)$. Therefore, Eq. (2) can be viewed as Bayesian inference with a prior distribution $p_{\mathrm{base}}(x)$ and the likelihood $\exp\left(r(x)/{\tau}\right)$ [26, 25]. In practice, the pre-trained model not only provides the prior distribution $p_{\mathrm{base}}(x)$ but also provides parameter initialization for the policy network. Figure 1(b) illustrates the RL fine-tuning process. The variational inference provides an alternative to the Markov chain Monte Carlo (MCMC) sampling carried out in Ref. [1]. Here, larger $\tau$ corresponds to higher temperature and hence stronger entropy regularization. Such optimization is also equivalent to the variational free energy calculations carried out for statistical mechanics problems [27, 28], where the two terms in the objective function Eq. (1) correspond to expected energy and entropy, respectively. We employ the proximal policy optimization (PPO) [29] algorithm to maximize the expected reward. The implementation details are provided in Appendix C.
@@
The CrystalFormer is an autoregressive transformer which models the tokenized representation of crystal structures with explicit knowledge of space group symmetries. The representation includes the space group number, Wyckoff letter, chemical element, and fractional coordinates of each symmetry inequivalent atom, and finally, the lattice parameters [1]. For example, the crystal LaH10 in the $Fm\bar{3}m$ (No. 225) space group with the cubic conventional unit cell of the length $5.1\AA$ is represented as the sequence ”225-a-La-0-0-0-c-H-1/4-1/4-1/4-f-H-0.375-0.375-0.375-5.1-5.1-5.1-90-90-90”, which contains all the necessary information of the crystal. Key to the design of CrystalFormer is the sequential nature of Wyckoff letters which decrease in site symmetry in an alphabet order. In this regard, the CrystalFormer leverages Nature’s codebook–the Wyckoff position table–for a multimodal probabilistic modeling. Although the representation appears to lack a sense of spatial geometry, unsupervised pre-training allows the model to ingest solid-state chemistry knowledge by compressing the crystalline material database into the model parameters. Furthermore, the RL phase may allow it to extract the geometry and physics from MLIP and property prediction models trained with supervised approaches.
@@ III.1 Reinforcement learning from MLIP for improved stability of generated materials
As the first application of the reinforcement fine-tuning of CrystalFormer, we choose the energy above the convex hull as the reward signal, i.e., $r(x)=-E_{\mathrm{hull}}(x)$. The energy above the convex hull is the key metric for assessing the stability of a material, with lower values indicating greater stability. To predict the energy and calculate the energy above the convex hull based on the Alexandria convex hull [37], we utilize the Orb model [13]. This choice allows us to obtain accurate energy predictions and assess the stability of materials relative to the convex hull fast. It is important to emphasize that our approach is not limited to the Orb model; any universal MLIP that can compute energy can serve as the reward model. More accurate and robust MLIP yields better reward signals, enhancing alignment with DFT calculations. Benchmarks on universal MLIP such as [38] provide an interactive leaderboard to select the favorable one for the fine-tuning purpose. We evaluate the performance by measuring the fraction of structures that are stable, unique, and novel (S.U.N.) [4]. The stability is defined as the energy above the convex hull being less than 0.1 eV/atom, uniqueness is defined as the generated structure not being identical to any other generated structures, and novelty is defined as the generated structure not being present in the training dataset.
@@
Figure 2(a) shows the expected energy above the convex hull of generated samples, which shows steady improvement during the training process. Figure 2(b) shows the KL divergence between the policy and the pre-trained base model $\mathrm{KL}(p_{\theta}\|p_{\mathrm{base}})=\mathop{\mathbb{E}}_{x\sim p_{\theta}(x)}\left[\ln\left(\frac{p_{\theta}(x)}{p_{\mathrm{base}}(x)}\right)\right]$, which is orders of magnitude smaller than change in the negative log-likelihood during pre-training [1].
@@
Figure 3: The histogram of energy above convex hull for the crystal samples from the pre-trained base model and the ones relaxed by the Orb model [13]. In comparison, reinforcement fine-tuned of the model significantly reduces the energy above convex hull of generated materials. The red dashed line indicates the threshold of 0.1 eV/atom for stable materials. Relaxation ratio from 44.7% in the base model to 57.8%, while RL fine-tuning improves the ratio of stable materials to 73.4% even without relaxation.
Figure 4: Fraction of generated structures that are stable, novel and unique (S.U.N.) for 14 space groups spanning the seven crystal systems. For each space group, two side-by-side bars correspond to the samples from the pre-trained (“Base”) and RL fine-tuned (“RL”) models. The S.U.N. structures form a subset of stable materials each space group. The shaded regions separate different crystal systems.
@@
Figure 3 shows the histogram of the energy above the convex hull for the pre-trained model and the reinforcement fine-tuned model. For each model, 1000 structures were sampled, with their space groups drawn from the training data distribution. For the pre-trained model, 44.7% of the generated structures show $E_{\mathrm{hull}}<0.1$ eV/atom. Relaxation improves this ratio to 57.8%. On the other hand, reinforcement fine-tuning with respect to the energy above the convex hull enhances the ratio to 73.4%. Such substantial improvement is due to that the reinforcement fine-tuning reshapes the distribution with a global modification of the Wyckoff sites, fractional coordinates, and lattice parameters while relaxation only modifies the materials locally. Besides reducing the energy above the convex hull, the reinforcement fine-tuning also improves the structure validity metrics compared to the base model [2], see Appendix B.
@@
Figure 4 shows the fraction of generated structures that are stable and are also S.U.N. for 14 space groups spanning the seven crystal systems. We investigate the performance of the model on the same 14 space groups as Ref. [4]. For each space group, we generate 1000 structures and evaluate the quality of samples using the Orb model [13]. The results demonstrate that the reinforcement fine-tuned model generates a higher fraction of stable and S.U.N. structures in most cases, indicating that the RL method can enhance the model’s ability to generate S.U.N. structures even though we only award for the stability. Reinforcement fine-tuning with respect to the energy above the convex improves the stability across all space group (as it should). However, the improvement of the S.U.N. is impaired by the novelty criteria as there is strong constraints to the structural motifs in these high-symmetry systems. From the figure one also observes that the performance of CrystalFormer is worst for the least symmetric $P1$ space group where there is minimal symmetry-related inductive bias to be explored. This is in sharp contrast to Ref. [4] which shows the best S.U.N. in the $P1$ space group, highlighting the distinctive advantage of our method. In fact, for space group with higher symmetry which are more relevant to the natural distribution of inorganic materials [39], we observed that the performance of CrystalFormer is on par or even better than the ones shown in Fig. D8 of Ref. [4]. Lastly, while reinforcement fine-tuning does not affect the generation speed of the model, it increases the rate of generating S.U.N. samples from 3.46 to 4.89 per second.
@@
Figure 5 breaks down the performance enhancement in terms of the S.U.N. ratio coming from enlarging the pre-training dataset and different ways of fine-tuning. For these evaluations, we sample space group following the same distribution of the training dataset and use it as precondition to the CrystalFormer. First of all, one sees that enlarging the pre-training dataset from MP-20 to Alex-20 yields a significant improvement in the S.U.N. ratio, rising from 2.6% to 15.3%. Next, we also tested the improvement brought by supervised fine-tuning (SFT) on the pre-trained model. For SFT, we generate 10,000 crystal samples and relax them using the Orb models. Only those stable samples with $E_{\mathrm{hull}}<0.1$ eV/atom were used to finetune the pre-trained model. Figure 5 shows that SFT yields some improvement over the pre-trained model, further increasing the S.U.N. ratio from 15.3% to 18.3%. However, it is the reinforcement fine-tuned model that delivers the most significant increase in the S.U.N. ratio, leading to a final S.U.N. ratio of 21.6%.
@@ III.2 Reinforcement learning from the figure of merits for property-guided material generation
Figure 6: (a) The distribution of band gap $E_{g}$ and electronic dielectric constant $\varepsilon_{\mathrm{elec}}$ of material samples generated in the $Fm\bar{3}m$ space group (No.225). The two properties are predicted by the MEGNet and AnisoNet models, respectively. The dashed line indicates a constant FoM at values $\varepsilon_{\mathrm{elec}}\cdot E_{g}=10$eV. The inset shows average FoM of generated materials versus training steps. Two stars indicate the S.U.N. materials. (b) The histogram of energy above the convex hull predicted by the Orb model [13] for the crystal samples from the pre-trained and the RL fine-tuned model in $Fm\bar{3}m$ space group (No.225), respectively. The red dashed line indicates the threshold of 0.1 eV/atom and the $x$-axis is up to 0.4 eV/atom. RL improves the FoM of the generated samples towards high FoM dielectric materials at the cost of only slightly affecting the stability. (c) Two examples of discovered high FoM materials along with their properties verified by DFT calculation. We also show the tokenized material string generated with CrystalFormer (fractional coordinate and lattice parameters omitted).
Next, we apply the RL method for dielectric material discovery. The interplay between the electronic dielectric constant and the band gap reflects a fundamental constraint in condensed matter systems. Recent works have established that the electronic polarizability originates from virtual interband transitions governed by the quantum metric and inversely scales with the spectral gap [40], while the gap itself is bounded by the dielectric response and electron density [41]. Pursuing materials that simultaneously exhibit a large electronic dielectric constant and a wide band gap thus represents a key challenge in realizing highly polarizable yet robust insulators, where quantum geometry and energetic stability reach a delicate balance. We utilize the pre-trained MEGNet [15, 42] and AnisoNet [43] models to predict the band gap $E_{g}$ and electronic dielectric constant $\varepsilon_{\mathrm{elec}}$ of the generated samples, respectively. MEGNet achieves a mean absolute error (MAE) of 0.314 eV for band gap prediction on the Materials Project dataset [44], while AnisoNet achieves a MAE of 0.297 for electronic dielectric constant prediction on the Materials Project dielectric dataset [44, 45]. It is worth noting both models are trained on the data computed using the Perdew-Burke-Ernzerhof functional [46], which is known to underestimate the band gap. These data also omit corrections such as spin-orbit coupling or excitonic effects, which may be important for certain materials. More accurate band gap prediction model would require training on higher-fidelity datasets, such as those computed using hybrid functionals [47, 48], the GW method [49] or GW-BSE approach [50]. The electronic dielectric constant is a dimensionless number, which is the ratio of the permittivity of a substance to the permittivity of vacuum. The reward function is defined as the figure of merit (FoM), $r(x)=\varepsilon_{\mathrm{elec}}\cdot E_{g}$, which favors materials with both sizeable band gap and permittivity.
@@
Figure 6(a) shows the ML predicted band gap $E_{g}$ versus electronic dielectric constant $\varepsilon_{\mathrm{elec}}$ for generated samples in the $Fm\bar{3}m$ space group (No.225) using the base and fine-tuned model respectively. The reinforcement fine-tuning significantly improves the FoM, with the majority of generated samples exhibiting higher FoM values compared to the pre-trained base model. We observe that the RL method primarily enhances the FoM by increasing the band gap, rather than the dielectric constant. This outcome may be attributed to the limited and relatively low dielectric constant data used to train the dielectric model, potentially causing the model to predict lower dielectric constants. To address this issue, a more robust dielectric prediction model that captures a wider range of dielectric constants could be employed [51]. In contrast, the band gap prediction model demonstrates greater accuracy, resulting in the RL method favoring an increase in the band gap to optimize the FoM.
@@
We first generated 1000 samples and then use Orb model to filter out the unstable materials, followed by DFT verification of the top 50 candidates with the highest FoM in the remaining subset. More computational details are provided in Appendix D, and materials with $E_{\mathrm{hull}}<0.1$ eV/atom are listed in Table LABEL:table:dielectric_samples. We observed that many of these samples are fluoride, probably due to fluorine’s strong electronegativity increases the ionic character of bonds, so that to ensure a sizeable band gap. Moreover, many of these high FoM samples are composed of Cs elements, which may be due to the relatively large ionic radius of Cesium, leading to higher polarizability and thus an increased total dielectric constant of the material. Figure 6(b) shows the energy above the convex hull for the materials sampled from the base and fine-tuned model respectively, which suggests that the reinforcement fine-tuning does not significantly affect the stability of generated materials. Two examples of high FoM materials discovered by CrystalFormer-RL, Cs2LiLuF6 and CsCaF3, are shown in Fig 6(c). Cs2LiLuF6 belongs to the double perovskite [52] structure prototype. CsCaF3 is obtained by substituting Ba with Ca in CsBaF3, which is included in the Alex-20 dataset, and it crystallizes in the halide perovskite [52, 53] structure prototype.